# A Comparison-Relationship-Surrogate Evolutionary Algorithm for Multi-Objective Optimization

### v1

@@ 4 Empirical Studies of Performance
Problem n NSGA-II ParEGO K-RVEA MOEA/D-EGO CPS-MOEA REMO CRSEA WFG1 16 1.44e+0 (2.41e-1) $\approx$ 1.32e+0 (4.70e-3) + 1.35e+0 (4.90e-2) + 1.38e+0 (2.90e-2) $\approx$ 1.36e+0 (1.33e-2) $\approx$ 1.30e+0 (5.23e-2) + 1.50e+0 (1.69e-1) 32 1.57e+0 (3.33e-1) $\approx$ 1.32e+0 (3.71e-3) + 1.33e+0 (3.50e-2) + 1.40e+0 (4.26e-2) + 1.37e+0 (1.22e-2) + 1.32e+0 (4.91e-2) + 1.53e+0 (1.79e-1) 64 1.49e+0 (2.52e-1) + - 1.36e+0 (4.86e-2) + 1.42e+0 (4.65e-2) + 1.37e+0 (8.70e-3) + 1.34e+0 (4.97e-2) + 1.58e+0 (1.73e-1) WFG2 16 5.14e-1 (1.31e-1) + 6.32e-1 (3.14e-2) - 4.85e-1 (6.56e-2) + 5.37e-1 (4.14e-2) + 5.72e-1 (5.51e-2) $\approx$ 4.87e-1 (6.19e-2) + 5.66e-1 (6.39e-2) 32 6.34e-1 (1.53e-1) + 6.69e-1 (2.12e-2) $\approx$ 6.56e-1 (2.73e-2) $\approx$ 5.97e-1 (4.17e-2) + 6.05e-1 (5.70e-2) + 5.43e-1 (6.10e-2) + 6.58e-1 (4.87e-2) 64 6.51e-1 (1.22e-1) + - 7.16e-1 (4.09e-2) $\approx$ 6.68e-1 (3.26e-2) + 6.54e-1 (4.50e-2) + 6.15e-1 (5.01e-2) + 7.09e-1 (2.78e-2) WFG3 16 4.00e-1 (5.11e-2) - 6.04e-1 (2.17e-2) - 3.04e-1 (6.78e-2) - 4.90e-1 (2.61e-2) - 5.21e-1 (3.37e-2) - 4.27e-1 (4.36e-2) - 2.24e-1 (4.96e-2) 32 4.93e-1 (4.00e-2) - 6.54e-1 (1.53e-2) - 5.35e-1 (4.02e-2) - 5.35e-1 (2.31e-2) - 5.76e-1 (2.55e-2) - 5.24e-1 (3.69e-2) - 4.61e-1 (4.58e-2) 64 5.86e-1 (2.48e-2) + - 6.33e-1 (1.61e-2) - 6.38e-1 (1.49e-2) - 6.16e-1 (2.79e-2) $\approx$ 5.96e-1 (2.15e-2) + 6.15e-1 (2.38e-2) WFG4 16 2.68e-1 (7.70e-2) - 3.86e-1 (2.41e-2) - 3.06e-1 (2.78e-2) - 3.36e-1 (6.10e-2) - 3.09e-1 (2.73e-2) - 2.66e-1 (2.94e-2) - 1.49e-1 (1.45e-2) 32 3.36e-1 (6.65e-2) - 4.75e-1 (1.97e-2) - 3.78e-1 (2.01e-2) - 3.85e-1 (5.12e-2) - 3.63e-1 (1.94e-2) - 3.24e-1 (4.93e-2) - 2.40e-1 (2.49e-2) 64 4.01e-1 (9.31e-2) $\approx$ - 4.21e-1 (2.72e-2) - 4.00e-1 (2.09e-2) - 3.88e-1 (2.22e-2) $\approx$ 3.76e-1 (3.46e-2) $\approx$ 3.83e-1 (2.56e-2) WFG5 16 3.66e-1 (4.02e-2) - 1.09e-1 (1.38e-2) - 1.78e-1 (2.93e-2) - 2.20e-1 (4.24e-2) - 3.30e-1 (2.44e-2) - 2.84e-1 (4.19e-2) - 9.37e-2 (1.50e-2) 32 4.83e-1 (5.60e-2) - 1.66e-1 (5.15e-2) - 3.39e-1 (3.24e-2) - 3.09e-1 (5.78e-2) - 3.78e-1 (2.42e-2) - 4.03e-1 (3.84e-2) - 1.17e-1 (1.48e-2) 64 5.69e-1 (4.75e-2) - - 4.88e-1 (3.11e-2) - 4.13e-1 (4.95e-2) - 4.20e-1 (1.95e-2) - 5.13e-1 (3.44e-2) - 1.79e-1 (2.58e-2) WFG6 16 4.11e-1 (5.76e-2) - 2.96e-1 (6.06e-2) $\approx$ 5.04e-1 (5.19e-2) - 4.45e-1 (8.56e-2) - 6.24e-1 (3.75e-2) - 4.75e-1 (5.08e-2) - 3.26e-1 (8.07e-2) 32 5.58e-1 (4.90e-2) - 2.84e-1 (6.22e-2) $\approx$ 5.81e-1 (4.22e-2) - 6.20e-1 (8.78e-2) - 7.23e-1 (3.89e-2) - 5.72e-1 (4.80e-2) - 3.46e-1 (1.62e-1) 64 6.75e-1 (3.76e-2) - - 6.93e-1 (2.55e-2) - 6.89e-1 (5.92e-2) - 7.51e-1 (2.78e-2) - 6.72e-1 (3.18e-2) - 4.66e-1 (1.03e-1) WFG7 16 3.62e-1 (5.56e-2) - 4.69e-1 (5.63e-2) - 4.23e-1 (3.19e-2) - 4.37e-1 (2.62e-2) - 4.32e-1 (2.24e-2) - 3.64e-1 (5.33e-2) - 1.72e-1 (2.94e-2) 32 4.51e-1 (5.29e-2) - 5.83e-1 (2.44e-2) - 5.28e-1 (1.77e-2) - 4.93e-1 (3.09e-2) - 4.85e-1 (1.74e-2) - 4.48e-1 (3.01e-2) - 3.53e-1 (3.03e-2) 64 5.31e-1 (5.64e-2) - - 5.76e-1 (9.12e-3) - 5.43e-1 (1.65e-2) - 5.18e-1 (2.08e-2) - 5.13e-1 (2.97e-2) $\approx$ 5.05e-1 (2.14e-2) WFG8 16 4.75e-1 (4.88e-2) - 6.46e-1 (4.49e-2) - 4.32e-1 (3.53e-2) - 5.79e-1 (2.84e-2) - 5.74e-1 (4.10e-2) - 4.65e-1 (4.16e-2) - 3.50e-1 (3.54e-2) 32 5.10e-1 (5.30e-2) - 6.73e-1 (2.88e-2) - 4.71e-1 (3.25e-2) - 5.75e-1 (2.38e-2) - 5.83e-1 (2.93e-2) - 4.87e-1 (4.13e-2) - 4.31e-1 (3.41e-2) 64 5.68e-1 (5.70e-2) - - 5.85e-1 (2.68e-2) - 6.04e-1 (2.51e-2) - 5.93e-1 (2.85e-2) - 5.17e-1 (2.93e-2) - 4.86e-1 (3.17e-2) WFG9 16 3.90e-1 (7.57e-2) - 3.61e-1 (6.84e-2) - 3.84e-1 (8.81e-2) - 5.27e-1 (6.49e-2) - 3.96e-1 (5.62e-2) - 3.66e-1 (1.10e-1) - 2.13e-1 (4.14e-2) 32 5.25e-1 (1.01e-1) - 4.17e-1 (7.48e-2) $\approx$ 6.23e-1 (1.02e-1) - 6.48e-1 (4.87e-2) - 4.46e-1 (4.53e-2) - 5.49e-1 (1.29e-1) - 3.87e-1 (1.11e-1) 64 6.59e-1 (6.99e-2) $\approx$ - 7.77e-1 (3.87e-2) - 7.30e-1 (4.86e-2) - 5.48e-1 (8.14e-2) + 6.83e-1 (8.90e-2) $\approx$ 6.74e-1 (8.00e-2) +/-$\approx$ 5/18/4 2/12/4 4/21/2 5/21/1 5/18/4 7/17/3
@@
Parameters for the new comparison-relationship-surrogate optimizer were selected as follows. The number of new function evaluations ($\mu$ in Alg. 1) was set to two and the number of “model generations” where the optimizer is run only on the surrogate model ($w_{\text{max}}$) was set to 32. Two hidden layers of size 64 were used in the feature extraction networks and one hidden layer of size 32 was used for the comparison networks. The number of epochs in the initial training of the model was 128 with 16 epochs used in every retraining in further generations. The learning rate was 0.001 at the beginning of each training cycle. An L2 regularization loss term was used (ie through the AdamW optimizer [42]) with a weight of $\lambda=0.1$.
@@ 5 Performance on a Real-World Problem
As another benchmark of the comparison-relationship-surrogate optimizer, we chose to evaluate it on a practical medium-scale problem from the field of accelerator physics. This problem is a photoinjector optimization for ultrafast electron diffraction applications [52] with 13 decision variables. Due to the high computational cost of the test, it is restricted to a comparison between the new comparison-based evolutionary algorithm and a single conventional optimizer (NSGA-II here).
@@ 6 Conclusion
In this work, we investigated the use of a comparison-relationship-surrogate model to improve the quality of solutions achievable with multi-objective genetic algorithms on a fixed budget of function evaluations. Unlike many previous SAEAs which predict the objective functions directly or the domination relationships between individuals, this method is inherently invariant to monotonic transformations of the objective functions and forms a naturally balanced classification problem.
We showed how a model that only has access to information about the comparison relationships between the objective functions of candidate solutions can be used to drive a basic non-dominated sorting evolutionary algorithm. This algorithm can quickly find promising solutions from the model without evaluating the expensive objective functions. We then demonstrated how the model can be periodically updated by using the real objective functions to create new training data from the candidate solutions after some maximum number of “model generations” has been reached. The model’s accuracy will be increased with each new set of evaluated solutions and the process can be repeated until the budget of function evaluations is exhausted.
Our numerical experiments found that the comparison-relationship-surrogate optimizer produced among the best solutions of all the algorithms studied for many medium-scale, biobjective problems (specifically WFG3-WFG9 in Tab. 2). Tests on a real-world optimization problem also showed the ability of the algorithm to find a larger solution hypervolume than the tested conventional genetic optimizer (NSGA-II).
The new algorithm had difficulty preserving the quality of solutions when the number of objectives grew during tests. This can be seen from its results on the DTLZ problem set (Tab. 1). One issue is that the performance of domination predictions from the comparison-relationship model degrades as the number of dimensions grows large. An accurate prediction of an individual that dominates another requires that all of the underlying comparisons were predicted correctly which becomes more difficult as objectives are added. Finding ways to avoid this fall-off in recall with a growing number of objectives could be a path toward improved performance.
Another challenge facing the algorithm on many-objective problems was the difficulty of maintaining solution diversity. The comparison-relationship model is unique among the classification-based SAEAs in that it enables predicting the individuals with extreme values of the objective functions for some approximation of crowding distance sorting. However, the model doesn’t have information about the absolute position of individuals in objective space and so beyond that, many existing diversity preservation mechanisms cannot be used. Investigating methods of improving solution diversity with a comparison surrogate might be a valuable future direction of research. The addition of constraints using a model of feasibility may be another interesting extension to the algorithm as it can be naturally added to the prediction of domination in the same way as in Ref. [38].
The algorithm described in this paper represents a new type of surrogate-assisted evolutionary algorithm in the growing genre of classification-based efficient multi-objective optimizers. These types of surrogates promise improved scaling with the number of decision variables as well as symmetry under monotonic transformation of the objective functions which is also reflected in many evolutionary algorithms [34]. With continued work in this area, these algorithms could help to improve the results available to the scientific and industrial users of optimization algorithms that face expensive multi-objective problems on a limited budget.

### v2

@@ 4 Empirical Studies of Performance
Problem n NSGA-II ParEGO K-RVEA MOEA/D-EGO CPS-MOEA REMO PC-SAEA CRSEA WFG1 16 1.44e+0 (2.41e-1) $\approx$ 1.32e+0 (4.70e-3) + 1.35e+0 (4.90e-2) + 1.38e+0 (2.90e-2) $\approx$ 1.36e+0 (1.33e-2) $\approx$ 1.30e+0 (5.23e-2) + 1.42e+0 (1.27e-1) $\approx$ 1.50e+0 (1.69e-1) 32 1.57e+0 (3.33e-1) $\approx$ 1.32e+0 (3.71e-3) + 1.33e+0 (3.50e-2) + 1.40e+0 (4.26e-2) + 1.37e+0 (1.22e-2) + 1.32e+0 (4.91e-2) + 1.58e+0 (2.94e-1) $\approx$ 1.53e+0 (1.79e-1) 64 1.49e+0 (2.52e-1) + - 1.36e+0 (4.86e-2) + 1.42e+0 (4.65e-2) + 1.37e+0 (8.70e-3) + 1.34e+0 (4.97e-2) + 1.40e+0 (1.21e-1) + 1.58e+0 (1.73e-1) WFG2 16 5.14e-1 (1.31e-1) + 6.32e-1 (3.14e-2) - 4.85e-1 (6.56e-2) + 5.37e-1 (4.14e-2) + 5.72e-1 (5.51e-2) $\approx$ 4.87e-1 (6.19e-2) + 5.47e-1 (6.66e-2) $\approx$ 5.66e-1 (6.39e-2) 32 6.34e-1 (1.53e-1) + 6.69e-1 (2.12e-2) $\approx$ 6.56e-1 (2.73e-2) $\approx$ 5.97e-1 (4.17e-2) + 6.05e-1 (5.70e-2) + 5.43e-1 (6.10e-2) + 6.64e-1 (4.92e-2) $\approx$ 6.58e-1 (4.87e-2) 64 6.51e-1 (1.22e-1) + - 7.16e-1 (4.09e-2) $\approx$ 6.68e-1 (3.26e-2) + 6.54e-1 (4.50e-2) + 6.15e-1 (5.01e-2) + 6.97e-1 (4.31e-2) $\approx$ 7.09e-1 (2.78e-2) WFG3 16 4.00e-1 (5.11e-2) - 6.04e-1 (2.17e-2) - 3.04e-1 (6.78e-2) - 4.90e-1 (2.61e-2) - 5.21e-1 (3.37e-2) - 4.27e-1 (4.36e-2) - 4.31e-1 (4.07e-2) - 2.24e-1 (4.96e-2) 32 4.93e-1 (4.00e-2) - 6.54e-1 (1.53e-2) - 5.35e-1 (4.02e-2) - 5.35e-1 (2.31e-2) - 5.76e-1 (2.55e-2) - 5.24e-1 (3.69e-2) - 5.53e-1 (2.22e-2) - 4.61e-1 (4.58e-2) 64 5.86e-1 (2.48e-2) + - 6.33e-1 (1.61e-2) - 6.38e-1 (1.49e-2) - 6.16e-1 (2.79e-2) $\approx$ 5.96e-1 (2.15e-2) + 6.31e-1 (9.93e-3) - 6.15e-1 (2.38e-2) WFG4 16 2.68e-1 (7.70e-2) - 3.86e-1 (2.41e-2) - 3.06e-1 (2.78e-2) - 3.36e-1 (6.10e-2) - 3.09e-1 (2.73e-2) - 2.66e-1 (2.94e-2) - 3.00e-1 (5.66e-2) - 1.49e-1 (1.45e-2) 32 3.36e-1 (6.65e-2) - 4.75e-1 (1.97e-2) - 3.78e-1 (2.01e-2) - 3.85e-1 (5.12e-2) - 3.63e-1 (1.94e-2) - 3.24e-1 (4.93e-2) - 3.45e-1 (3.48e-2) - 2.40e-1 (2.49e-2) 64 4.01e-1 (9.31e-2) $\approx$ - 4.21e-1 (2.72e-2) - 4.00e-1 (2.09e-2) - 3.88e-1 (2.22e-2) $\approx$ 3.76e-1 (3.46e-2) $\approx$ 3.98e-1 (2.38e-2) - 3.83e-1 (2.56e-2) WFG5 16 3.66e-1 (4.02e-2) - 1.09e-1 (1.38e-2) - 1.78e-1 (2.93e-2) - 2.20e-1 (4.24e-2) - 3.30e-1 (2.44e-2) - 2.84e-1 (4.19e-2) - 4.37e-1 (3.56e-2) - 9.37e-2 (1.50e-2) 32 4.83e-1 (5.60e-2) - 1.66e-1 (5.15e-2) - 3.39e-1 (3.24e-2) - 3.09e-1 (5.78e-2) - 3.78e-1 (2.42e-2) - 4.03e-1 (3.84e-2) - 4.98e-1 (1.95e-2) - 1.17e-1 (1.48e-2) 64 5.69e-1 (4.75e-2) - - 4.88e-1 (3.11e-2) - 4.13e-1 (4.95e-2) - 4.20e-1 (1.95e-2) - 5.13e-1 (3.44e-2) - 5.78e-1 (1.89e-2) - 1.79e-1 (2.58e-2) WFG6 16 4.11e-1 (5.76e-2) - 2.96e-1 (6.06e-2) $\approx$ 5.04e-1 (5.19e-2) - 4.45e-1 (8.56e-2) - 6.24e-1 (3.75e-2) - 4.75e-1 (5.08e-2) - 5.37e-1 (3.64e-2) - 3.26e-1 (8.07e-2) 32 5.58e-1 (4.90e-2) - 2.84e-1 (6.22e-2) $\approx$ 5.81e-1 (4.22e-2) - 6.20e-1 (8.78e-2) - 7.23e-1 (3.89e-2) - 5.72e-1 (4.80e-2) - 6.42e-1 (3.11e-2) - 3.46e-1 (1.62e-1) 64 6.75e-1 (3.76e-2) - - 6.93e-1 (2.55e-2) - 6.89e-1 (5.92e-2) - 7.51e-1 (2.78e-2) - 6.72e-1 (3.18e-2) - 7.15e-1 (3.99e-2) - 4.66e-1 (1.03e-1) WFG7 16 3.62e-1 (5.56e-2) - 4.69e-1 (5.63e-2) - 4.23e-1 (3.19e-2) - 4.37e-1 (2.62e-2) - 4.32e-1 (2.24e-2) - 3.64e-1 (5.33e-2) - 3.99e-1 (3.77e-2) - 1.72e-1 (2.94e-2) 32 4.51e-1 (5.29e-2) - 5.83e-1 (2.44e-2) - 5.28e-1 (1.77e-2) - 4.93e-1 (3.09e-2) - 4.85e-1 (1.74e-2) - 4.48e-1 (3.01e-2) - 4.71e-1 (2.74e-2) - 3.53e-1 (3.03e-2) 64 5.31e-1 (5.64e-2) - - 5.76e-1 (9.12e-3) - 5.43e-1 (1.65e-2) - 5.18e-1 (2.08e-2) - 5.13e-1 (2.97e-2) $\approx$ 5.32e-1 (2.52e-2) - 5.05e-1 (2.14e-2) WFG8 16 4.75e-1 (4.88e-2) - 6.46e-1 (4.49e-2) - 4.32e-1 (3.53e-2) - 5.79e-1 (2.84e-2) - 5.74e-1 (4.10e-2) - 4.65e-1 (4.16e-2) - 5.63e-1 (4.06e-2) - 3.50e-1 (3.54e-2) 32 5.10e-1 (5.30e-2) - 6.73e-1 (2.88e-2) - 4.71e-1 (3.25e-2) - 5.75e-1 (2.38e-2) - 5.83e-1 (2.93e-2) - 4.87e-1 (4.13e-2) - 5.71e-1 (4.74e-2) - 4.31e-1 (3.41e-2) 64 5.68e-1 (5.70e-2) - - 5.85e-1 (2.68e-2) - 6.04e-1 (2.51e-2) - 5.93e-1 (2.85e-2) - 5.17e-1 (2.93e-2) - 5.76e-1 (2.54e-2) - 4.86e-1 (3.17e-2) WFG9 16 3.90e-1 (7.57e-2) - 3.61e-1 (6.84e-2) - 3.84e-1 (8.81e-2) - 5.27e-1 (6.49e-2) - 3.96e-1 (5.62e-2) - 3.66e-1 (1.10e-1) - 4.18e-1 (7.64e-2) - 2.13e-1 (4.14e-2) 32 5.25e-1 (1.01e-1) - 4.17e-1 (7.48e-2) $\approx$ 6.23e-1 (1.02e-1) - 6.48e-1 (4.87e-2) - 4.46e-1 (4.53e-2) - 5.49e-1 (1.29e-1) - 6.15e-1 (6.07e-2) - 3.87e-1 (1.11e-1) 64 6.59e-1 (6.99e-2) $\approx$ - 7.77e-1 (3.87e-2) - 7.30e-1 (4.86e-2) - 5.48e-1 (8.14e-2) + 6.83e-1 (8.90e-2) $\approx$ 7.14e-1 (2.67e-2) - 6.74e-1 (8.00e-2) +/-$\approx$ 5/18/4 2/12/4 4/21/2 5/21/1 5/18/4 7/17/3 1/21/5
@@
Parameters for the new comparison-relationship-surrogate optimizer were selected as follows. The number of new function evaluations ($\mu$ in Alg. 1) was set to two and the number of “model generations” where the optimizer is run only on the surrogate model ($w_{\text{max}}$) was set to 32. We discuss more about this choice in a sensitivity study on the parameter in the following subsection. Two hidden layers of size 64 were used in the feature extraction networks and one hidden layer of size 32 was used for the comparison networks. The number of epochs in the initial training of the model was 128 with 16 epochs used in every retraining in further generations. The learning rate was 0.001 at the beginning of each training cycle. An L2 regularization loss term was used (i.e. through the AdamW optimizer [46]) with a weight of $\lambda=0.1$.
@@
4.1 Sensitivity to Number of Iterations Using Model
In this section, we evaluate the sensitivity of the CRSEA algorithm to the number of iterations of the evolutionary algorithm run using the model only ($w_{\text{max}}$ in Alg. 1). This parameter should be tuned to squeeze the most information out of the model without drifting too far into the regions where it does not have training data. For a test of how this parameter affects the performance of the algorithm, we have rerun CRSEA on the WFG problem suite with $n=16$ as we vary $w_{\text{max}}$. We report this data in Tab. 3. All of the settings besides $w_{\text{max}}$ were kept the same as in previous experiments. Every problem was evaluated 32 times and the Wilcoxon rank sum test used to compare them to the suggested set-point. These results suggest $w_{\text{max}}=32$ as a reasonable choice for this set of problems.
Table 3: Sensitivity study of the $w_{\text{max}}$ parameter in CRSEA. Statistical results (mean with standard deviation in parenthesis) are shown for the value of the IGD metric of 32 runs using each setting on the WFG (n=16) problems. Cells which are among the best performing for that problem up to a 95% significance level are highlighted.
Problem n CRSEA ($w_{\text{max}}=4$) CRSEA ($w_{\text{max}}=8$) CRSEA ($w_{\text{max}}=16$) CRSEA ($w_{\text{max}}=32$) CRSEA ($w_{\text{max}}=64$) WFG1 16 1.84e+0 (4.43e-2) - 1.75e+0 (7.98e-2) - 1.65e+0 (1.54e-1) - 1.47e+0 (1.86e-1) 1.42e+0 (1.43e-1) $\approx$ WFG2 16 5.46e-1 (6.70e-2) $\approx$ 5.48e-1 (6.01e-2) $\approx$ 5.63e-1 (6.22e-2) $\approx$ 5.57e-1 (5.33e-2) 5.90e-1 (5.35e-2) - WFG3 16 3.08e-1 (6.37e-2) - 2.57e-1 (4.37e-2) - 2.44e-1 (4.04e-2) $\approx$ 2.32e-1 (4.90e-2) 2.15e-1 (4.22e-2) $\approx$ WFG4 16 1.99e-1 (3.57e-2) - 1.78e-1 (2.89e-2) - 1.62e-1 (2.17e-2) - 1.50e-1 (1.37e-2) 1.52e-1 (1.30e-2) $\approx$ WFG5 16 1.81e-1 (2.99e-2) - 1.13e-1 (2.15e-2) - 9.52e-2 (7.45e-3) - 9.15e-2 (9.40e-3) 9.04e-2 (6.84e-3) $\approx$ WFG6 16 4.14e-1 (6.30e-2) - 3.34e-1 (6.17e-2) $\approx$ 3.24e-1 (6.09e-2) $\approx$ 3.16e-1 (7.45e-2) 2.80e-1 (6.75e-2) + WFG7 16 2.53e-1 (4.37e-2) - 2.03e-1 (4.59e-2) - 1.88e-1 (3.82e-2) $\approx$ 1.71e-1 (3.28e-2) 1.76e-1 (3.86e-2) $\approx$ WFG8 16 3.91e-1 (3.02e-2) - 3.65e-1 (3.46e-2) - 3.61e-1 (3.65e-2) - 3.47e-1 (3.06e-2) 3.51e-1 (3.57e-2) $\approx$ WFG9 16 2.92e-1 (9.69e-2) - 2.46e-1 (6.59e-2) - 2.24e-1 (5.34e-2) $\approx$ 2.07e-1 (6.07e-2) 2.25e-1 (4.68e-2) $\approx$ +/-$\approx$ 0/8/1 0/7/2 0/4/5 1/1/7
@@ 5 Performance on a Real-World Problem
As another benchmark of the comparison-relationship-surrogate optimizer, we chose to evaluate it on a practical medium-scale problem from the field of accelerator physics. This problem is a photoinjector optimization for ultrafast electron diffraction applications [54] with 13 decision variables. Due to the high computational cost of the test (20 to 30 minutes per function evaluation) and limits on the programming language the simulation tool and its interfaces are written in, it is restricted to a comparison between the new comparison-based evolutionary algorithm and a single conventional optimizer (NSGA-II here).
@@ 6 Conclusion
In this work, we investigated the use of a comparison-relationship-surrogate model to improve the quality of solutions achievable with multi-objective genetic algorithms on a fixed budget of function evaluations. This type of model, given two individuals, will predict which of them has the smaller value of each objective function. Domination relationships can then be inferred after a “cleaning” procedure allowing the model to be integrated into a non-dominated sorting genetic algorithm. Its performance was then compared to that of a selection of other algorithms on the DTLZ and WFG test suites as well as a real-world problem from the field of accelerator physics.
Our numerical experiments found that the comparison-relationship-surrogate optimizer produced among the best solutions of all the algorithms studied for many medium-scale, biobjective problems (specifically WFG3-WFG9 in Tab. 2). Tests on the real-world optimization problem also showed the ability of the algorithm to find a larger solution hypervolume than the tested conventional genetic optimizer (NSGA-II). Results were not as promising on the DTLZ test suite where solution diversity may have limited performance.
Finding diversity preservation methods which are compatible with predictions from the comparison-relationship-surrogate model will be an important line of research coming out of this work. Inclusion of constraints using a model of feasibility is also a natural extension to the work as it can be naturally added to the prediction of domination in the same way as in Ref. [42]. With continued work in this area, these algorithms could help to improve the results available to the scientific and industrial users of optimization algorithms that face expensive multi-objective problems on a limited budget.